# ul. Pasteura 5, 02-093 Warsaw, Poland

### v1

@@ Appendix A Details on the benchmark scenario
In this appendix we discuss the scenario leading to the GW spectrum reported in Fig. 5. We emphasize that the behaviour is non-generic and the scenario we are studying is somewhat tuned. We believe, however, that it illustrates certain interesting aspects of the physics of cosmic strings with varying tension. This case, in which the overall volume mode drives the variation of the tension, only gives a meaningful signal in this sort of tuned scenario, where the period of matter domination is naturally large (and so one must aim to get as many e-folds of kination as possible). However, other configurations where the string tension varies as a function of other moduli as those studied in Brunelli:2025ems could give less e-folds of matter domination, and we believe our discussion makes explicit the physics involved. We have also taken as a proxy the PTA signal but there is nothing preventing from studying other late-time volume configurations, which may lead to interesting signals with less tuning. These interesting avenues are left for future work.
The discussion is based on the potential reported in Cicoli:2024bwq (see Villa:2025zmj for a concise summary). The idea is to consider the supersymmetric gravity sector of low-energy string vacua (i.e: closed string moduli, their axionic counterparts and their fermionic superpartners) and couple it to a sufficiently sequestered source of supersymmetry breaking, such as an antibrane in a warped throat. This sort of system can be described in the nilpotent Goldstino formalism, started in Komargodski:2009rz (see Kallosh:2015nia and references therein for applications in string theory). For our present purposes, the outcome is that the most general perturbative scalar potential for the volume modulus compatible with other symmetries generic in string vacua (see Burgess:2020qsc ) is given by:
$V=\frac{1}{U}\left[\left(f^{\prime}W_{X}-3g^{\prime}W_{0}\right)^{2}\right.-\left.f^{\prime\prime}\left(fW_{X}^{2}-6gW_{X}W_{0}-9hW_{0}^{2}\right)\right]\,,$ (104)
@@
It is well known that the first contribution to $f^{\prime\prime}$ arises at order $f^{\prime\prime}\sim 1/\mathcal{V}^{3}$ in the form of the BBHL Becker:2002nn correction and dominates at small $\mathcal{V}$, providing the $\lambda=\sqrt{27/2}$ behaviour characteristic of LVS models. We neglect this contribution in our scenario (as could occur if the extra dimensions where a self-mirror CY manifold) and assume $f^{\prime\prime}=0$ at a sufficiently high order allowing for the behaviour $V\sim\mathcal{V}^{-8/3}$ at small $\tau$. The potential is then a perfect square. We assume on the grounds of effective field theory reasoning that quantum effects due to the antibrane generate a term of the form $g^{\prime}=A/\mathcal{V}^{a}$ (note that functions in string vacua generically admit expansions in inverse powers of cycle volumes). It is important to emphasize, however, that this correction is not the result of a stringy computation. Assuming it, the minimum is located at a position
$\mathcal{V}^{2/3}=\left(\frac{3AW_{0}}{W_{x}}\right)^{1/a}\,,$ (105)
@@
which can be exponentially large in the case where the Goldstino field is describing the effect of an antibrane in a warped throat. Indeed, in that case it is well known that $W_{x}^{2}\sim\exp\left(-8\pi K/3g_{s}M\right)$, with $K$ and $M$ a choice of integers (see Aparicio:2015psl ). We will take as a proxy $a=1$, which would occur for a logarithmic correction of the form $g(\mathcal{V})=A\log\mathcal{V}$, and so $A$ naturally carries powers of $g_{s}$. In terms of the canonically normalized field $\Phi$, we thus have:
$V=e^{-2\sqrt{2/3}\frac{\Phi}{M_{P}}}|W_{x}|^{2}/3-2Ae^{-\sqrt{6}\frac{\Phi}{M_{P}}}W_{x}W_{0}+3A^{2}W_{0}^{2}e^{-4\sqrt{\frac{2}{3}}\frac{\Phi}{M_{P}}}\,.$ (106)
@@
At small volume, the last term dominates providing the $\lambda=4\sqrt{\frac{2}{3}}$ behaviour. On a log-scale, this is a smaller slope than $\lambda=\sqrt{27/2}$ and so allows for more field range between the putative high-scale inflationary minimum, which we take at $\Lambda=10^{16}$ GeV, and the late-time vacuum. For our plots, we choose parameters $W_{x}=3AW_{0}(3/10^{10})^{2/3}$, ensuring a late-time volume $\mathcal{V}=10^{10}/3$, and $\sqrt{3}AW_{0}=10^{-5}$ which allows to have a large initial energy density $V(\Phi_{0})=10^{16}$ GeV with a moderate hierarchy $G\mu\simeq 8.8\times 10^{-4}$, although the initial volume $\mathcal{V}=7.5$ is unreasonably small ($\Phi=1.65M_{P}$). In addition, we take as initial conditions $\varepsilon_{0}=4.8\times 10^{-6}$, $x^{2}-\varepsilon_{0}=0.99$ and $y^{2}=0.01$, in the notation of Sec. 3.1.1. We stop the dynamical system evolution and patch to a period of early matter domination after the first oscillation around the minimum. Early matter domination ends at a time $1/\Gamma_{\Phi\to hh}$, with
$\Gamma_{\Phi\to hh}=\left(\frac{\mathcal{V}}{16\pi}\right)^{2}\frac{m_{\Phi}^{2}}{M_{P}^{3}}\,,$ (107)
@@
c.f. Eq. (45), and the mass of the volume is computed from Eq. (106), which for the choice of parameters indicated turns out to be $m_{\Phi}\simeq 8$ GeV.

### v2

@@ Appendix A Details on the benchmark scenario
In this appendix we discuss the scenario leading to the GW spectrum reported in Fig. 5. We emphasize that the behaviour is non-generic and the scenario we are studying is somewhat tuned. We believe, however, that it illustrates certain interesting aspects of the physics of cosmic strings with varying tension. This case, in which the overall volume mode drives the variation of the tension, only gives a meaningful signal in this sort of tuned scenario, where the period of matter domination is naturally large (and so one must aim to get as many e-folds of kination as possible). However, other configurations where the string tension varies as a function of other moduli as those studied in [80] could give less e-folds of matter domination, and we believe our discussion makes explicit the physics involved. We have also taken as a proxy the PTA signal but there is nothing preventing from studying other late-time volume configurations, which may lead to interesting signals with less tuning. These interesting avenues are left for future work.
The discussion is based on the potential reported in [39] (see [193] for a concise summary). The idea is to consider the supersymmetric gravity sector of low-energy string vacua (i.e: closed string moduli, their axionic counterparts and their fermionic superpartners) and couple it to a sufficiently sequestered source of supersymmetry breaking, such as an antibrane in a warped throat. This sort of system can be described in the nilpotent Goldstino formalism, started in [194] (see [195] and references therein for applications in string theory). For our present purposes, the outcome is that the most general perturbative scalar potential for the volume modulus compatible with other symmetries generic in string vacua (see [116]) is given by:
$V=\frac{1}{U}\left[\left(f^{\prime}W_{X}-3g^{\prime}W_{0}\right)^{2}\right.-\left.f^{\prime\prime}\left(fW_{X}^{2}-6gW_{X}W_{0}-9hW_{0}^{2}\right)\right]\,$, (A.1)
@@
It is well known that the first contribution to $f^{\prime\prime}$ arises at order $f^{\prime\prime}\sim 1/\mathcal{V}^{3}$ in the form of the BBHL [114] correction and dominates at small $\mathcal{V}$, providing the $\lambda=\sqrt{27/2}$ behaviour characteristic of LVS models. We neglect this contribution in our scenario (as could occur if the extra dimensions where a self-mirror CY manifold) and assume $f^{\prime\prime}=0$ at a sufficiently high order allowing for the behaviour $V\sim\mathcal{V}^{-8/3}$ at small $\tau$. The potential is then a perfect square. We assume on the grounds of effective field theory reasoning that quantum effects due to the antibrane generate a term of the form $g^{\prime}=\tilde{A}/\mathcal{V}^{a}$ (note that functions in string vacua generically admit expansions in inverse powers of cycle volumes). It is important to emphasize, however, that this correction is not the result of a stringy computation. Assuming it, the minimum is located at a position
$\mathcal{V}^{2/3}=\left(\frac{3\tilde{A}W_{0}}{W_{x}}\right)^{1/a}\,$, (A.2)
@@
which can be exponentially large in the case where the Goldstino field is describing the effect of an antibrane in a warped throat. Indeed, in that case it is well known that $W_{x}^{2}\sim\exp\left(-8\pi K/3g_{s}M\right)$, with $K$ and $M$ a choice of integers (see [196]). We will take as a proxy $a=1$, which would occur for a logarithmic correction of the form $g(\mathcal{V})=\tilde{A}\log\mathcal{V}$, and so $\tilde{A}$ naturally carries powers of $g_{s}$. In terms of the canonically normalized field $\Phi$, the perfect square contribution to the potential in Eq. (A.1) is: [^44] This contribution would then be uplifted to a de Sitter vacuum by a higher order correction to $f^{\prime\prime}$, which does not alter the discussion.
$V=e^{-2\sqrt{2/3}\frac{\Phi}{M_{P}}}|W_{x}|^{2}/3-2\tilde{A}e^{-\sqrt{6}\frac{\Phi}{M_{P}}}W_{x}W_{0}+3\tilde{A}^{2}W_{0}^{2}e^{-4\sqrt{\frac{2}{3}}\frac{\Phi}{M_{P}}}\,.$ (A.3)
This particular choice of parameters corresponds, in the notation of Eq. (3.11),to
$V_{0}=3\tilde{A}^{2}W_{0}^{2}\,,\qquad B=-2\tilde{A}W_{x}W_{0}/V_{0}\,,\qquad A=W_{x}^{2}/3V_{0}$
$\lambda_{3}=4\sqrt{2/3}\equiv\lambda\,,\qquad\lambda_{1}=\lambda/2\,,\qquad\lambda_{2}=3\lambda/4\,.$ (A.4)
@@
At small volume, the last term dominates providing the $\lambda=4\sqrt{\frac{2}{3}}$ behaviour in the notation of Eq. (3.11). On a log-scale, this is a smaller slope than $\lambda=\sqrt{27/2}$ and so allows for more field range between the putative high-scale inflationary minimum, which we take at $\Lambda=10^{16}$ GeV, and the late-time vacuum. For our plots, we choose parameters $W_{x}=3\tilde{A}W_{0}(3/10^{10})^{2/3}$, ensuring a late-time volume $\mathcal{V}=10^{10}/3$, and $\sqrt{3}\tilde{A}W_{0}=10^{-5}$ which allows to have a large initial energy density $V(\Phi_{0})=10^{16}$ GeV with a moderate hierarchy $G\mu\simeq 8.8\times 10^{-4}$, although the initial volume $\mathcal{V}=7.5$ is unreasonably small ($\Phi=1.65M_{P}$). In addition, we take as initial conditions $\varepsilon_{0}=4.8\times 10^{-6}$, $x^{2}-\varepsilon_{0}=0.99$ and $y^{2}=0.01$, in the notation of Sec. 3.1.1. We stop the dynamical system evolution and patch to a period of early matter domination after the first oscillation around the minimum. Early matter domination ends at a time $1/\Gamma_{\Phi\to hh}$, with
$\Gamma_{\Phi\to hh}=\left(\frac{\mathcal{V}}{16\pi}\right)^{2}\frac{m_{\Phi}^{2}}{M_{P}^{3}}\,$, (A.5)
@@
c.f. Eq. (3.24), and the mass of the volume is computed from Eq. (A.3), which for the choice of parameters indicated turns out to be $m_{\Phi}\simeq 8$ GeV.